A man sells two boats for $6,000.00 each. He made 20% profit on one boat and lost 20% on the other boat. Did he make money, lose money, or break even? How do you know?
Answer by mananth(16946)

First boat selling price $6000
profit 20%
His cost price = 6000*100/120 = $5000
-
Second boat loss = 20%
His cost price will be 6000*100/80 = $7500
Total of Sale = $12,000
Total cost = 5000+7500= $12500
You now know he lost $500